OpenStudy (mtalhahassan2):

Joe invests $1000 at the start of each year for five years and earns 6.3% simple interest on his investments. How much will all his investments be worth at the start of the fifth year

OpenStudy (mtalhahassan2):

@KamiBug

OpenStudy (solomonzelman):

at the start of the fifth year means that 4 years have passed.

OpenStudy (mtalhahassan2):

so basically for this question I am using the Arithmetric Series

OpenStudy (solomonzelman):

yes, pretty much

OpenStudy (solomonzelman):

you are adding the same simple interest every year

OpenStudy (mtalhahassan2):

ok so

OpenStudy (mtalhahassan2):

which formula is we using for solving this

OpenStudy (solomonzelman):

okay, the simple expression to calculate for just 4 years would be: \(\large\color{teal}{ 1000 \times 0.63 \times 4 }\) add this interest to the intial \(\large\color{slate}{ 1000 }\) you have, \(\large\color{teal}{ 1000+( 1000 \times 0.63 \times 4 ) }\)

OpenStudy (solomonzelman):

you can, if you like to, find the difference, which would be just the interest for a single year: \(\large\color{teal}{ d=1000 \times 0.63 }\)

OpenStudy (mtalhahassan2):

is we don't have to used any formula

OpenStudy (mtalhahassan2):